OpenStudy (anonymous):

this stupid thing is winning, help me beat it! The length of a rectangle is 5 more than twice the width. if the perimeter is 58cm, find the dimensions

OpenStudy (anonymous):

does it set up as 58 = 2+5?

OpenStudy (saifoo.khan):

no

OpenStudy (anonymous):

2x I mean

OpenStudy (anonymous):

width = 8 cm lenght = 21 cm

OpenStudy (anonymous):

whoa whoa, length is not 5w, it is 2w+5... right?

OpenStudy (anonymous):

5 more than twice the length is what it says

OpenStudy (anonymous):

my answer isn`t wrong Snapier?

OpenStudy (anonymous):

Lol. Complete misread. Yeah. So you have 2l+2w=58 l=2w+5 4w+20+2w=58 6w=48 w=8 Then l=2(8)+5 l=21

OpenStudy (anonymous):

Thats right^^

OpenStudy (anonymous):

I am just trying to make sure I understand how to set up the equation, I never checked yoru answer Paul

OpenStudy (anonymous):

ok, sweet... beers all around!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

So friend..... u have a system like that 5 + 2W = L .... A 2L + 2W = 58 .... B then, for example for solve it, take (B - A) this implie in a new equation C like that 2L - 5 = - L + 58 ... C then 3L = 63 \[\rightarrow\] L = 21 insert it in eq. A we have 5 + 2W = 21 solving \[\rightarrow 2W =16\] W = 8 Answer > \[L = 21 cm , W = 8 cm\]